OpenStudy (anonymous):

Find a cubic function with the given zeros. -2, 5, -6

OpenStudy (anonymous):

f(x) = x3 + 3x2 + 28x - 60 f(x) = x3 - 3x2 - 28x - 60 f(x) = x3 + 3x2 - 28x - 60 f(x) = x3 + 3x2 - 28x + 60

OpenStudy (anonymous):

(x+2)(x-5)(x+6) Expand that.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Multiply out; or FOIL.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Im stuck

OpenStudy (anonymous):

Stuck where?

OpenStudy (anonymous):

I dont know how to foil three sets of numbers

OpenStudy (anonymous):

Do two first. Then do the third one.

OpenStudy (anonymous):

so do (x+2)(x-5) and then (x-5)(x+6)

OpenStudy (anonymous):

No. Do (x+2)(x-5). For whatever you get; do that with (x+6)

OpenStudy (anonymous):

ohhhhhhhhhhhhhh thats soooo easy

OpenStudy (anonymous):

Let me see what you do.

OpenStudy (anonymous):

\[(x+2)\text{ }(x-5 )( x+6)=\left(x^2-3 x-10\right)( x+6) \]

OpenStudy (anonymous):

thats what I just got

OpenStudy (anonymous):

Go on from there. Just do the two as you would.